### v1

# Improving Coverage in Combined Prediction Sets with Weighted p-values

Abstract
Conformal prediction quantifies the uncertainty of machine learning models by augmenting point predictions with valid prediction sets, assuming exchangeability. For complex scenarios involving multiple trials, models, or data sources, conformal prediction sets can be aggregated to create a prediction set that captures the overall uncertainty, often improving precision. However, aggregating multiple prediction sets with individual $1-\alpha$ coverage inevitably weakens the overall guarantee, typically resulting in $1-2\alpha$ worst-case coverage. In this work, we propose a framework for the weighted aggregation of prediction sets, where weights are assigned to each prediction set based on their contribution. Our framework offers flexible control over how the sets are aggregated, achieving tighter coverage bounds that interpolate between the $1-2\alpha$ guarantee of the combined models and the $1-\alpha$ guarantee of an individual model depending on the distribution of weights. We extend our framework to data-dependent weights, and we derive a general procedure for data-dependent weight aggregation that maintains finite-sample validity. We demonstrate the effectiveness of our methods through experiments on synthetic and real data in the mixture-of-experts setting, and we show that aggregation with data-dependent weights provides a form of adaptive coverage.
1 Introduction
In recent years, machine learning models have achieved remarkable accuracy across a variety of predictive tasks [1, 2, 3]. Understanding the uncertainty associated with each prediction is essential to decision-making in real-world scenarios; however, the black box nature of many machine learning models hinders their deployment in safety-critical applications such as medical diagnosis [4, 5, 6, 7], industrial control systems [8, 9], and extreme weather forecasting [10, 11, 12]. Conformal prediction [13] has emerged as a popular wrapper method around machine learning models because it provides a statistically valid quantification of uncertainty. In particular, it transforms point predictions to prediction sets with finite-sample coverage guarantees, as long as the test data is exchangeable with the training data. Complex scenarios involving multiple predictions can therefore result in multiple conformal prediction sets—for example, when there are multiple trials, models, or data sources.
Aggregating multiple prediction sets provides a summary view of the uncertainty across predictions. Various aggregation methods exist, each with their own balance of sample efficiency and computational cost. For conformal prediction, where the popular split conformal variant [14] introduces a one-time random split, there are multiple methods of aggregating predictions to reduce the randomness over multiple splits: popular examples include cross-conformal [15], CV+ [16], and jackknife+ [16]. These aggregation methods are all symmetric, in that individual prediction sets contribute equally to the aggregate set. Although comparatively less studied, it is often useful to aggregate sets asymmetrically: that is, to weight sets based on their prior importance to the overall result [17]. Work in both symmetric and asymmetric aggregation establishes that aggregating individual prediction sets with $1-\alpha$ coverage guarantees results in an overall coverage guarantee of $1-2\alpha$ [18].
Our work is based on the observation that if the overall coverage guarantee for asymmetric aggregation reflects the contributions of the individual sets, then we can achieve a tighter guarantee than a constant $1-2\alpha$. Consider, as an extreme, the case where almost all the weight is on one set; then, because the coverage of that set dominates the aggregation, the overall coverage guarantee should be closer to the $1-\alpha$ guarantee of the dominant individual set. To this end, we propose a method for asymmetric aggregation where the overall coverage guarantee is based on the distribution of weights across the prediction sets. This approach leverages the results of Vovk and Wang [18] on averaging p-values to achieve tighter guarantees when there exists strong asymmetry in the importance of the prediction sets—i.e. when a single prediction set is significantly more important than the rest—with $1-2\alpha$ as the worst-case guarantee when the contributions of the prediction sets are more equal.
The results on averaging p-values require that the weights be independent of the prediction sets, and therefore of the data. However, it is often desirable for weights to reflect information about the data—such as in the popular mixture-of-experts setting [19], where expert weights are learned from data. To deal with this general case, we additionally propose a method for aggregating prediction sets with data-dependent weights via a linear transformation on the weighted average of p-values associated with each set. Our method allows us to construct a valid aggregate prediction set while also preserving the proportions of the weights assigned to the individual sets. Since data-dependent weights adjust the contributions of the individual sets based on the observed input, our method also achieves a form of conditional/data-adaptive coverage. We apply our method to the mixture-of-experts setting, and we demonstrate its effectiveness in experiments with real and synthetic data.
To summarize, our main contributions are as follows:
  • We propose a framework for the asymmetric aggregation of prediction sets based on weighted p-values (§4). With data-independent weights, our framework improves the coverage guarantee beyond the standard $1-2\alpha$, depending on the distribution of weights (Prop. 4.1).
  • We derive a general method for transforming a random variable to a p-value variable, which enables the construction of a prediction set with finite-sample guarantees (§5). Our method allows us to extend our framework to data-dependent weights, and we demonstrate that incorporating data-dependent weights provides a form of (locally) conditional coverage (Prop. 5.1, Prop. 5.2).
  • We apply our method to the mixture-of-experts setting, and we demonstrate that using a weighted aggregation of experts improves local validity (§6).
2 Background
We begin by reviewing preliminaries leading up to the basics of conformal prediction, with a focus on the widely-used split conformal method. To provide context for our method based on p-values, we establish the connection between the more typical quantile presentation of conformal prediction and its original p-value presentation [13].
Quantile and the empirical quantile function   Given a random variable $X\in\mathcal{X}$, the left $\alpha$-quantile of $X$, which we refer to as the $\alpha$-quantile of $X$ for simplicity, is defined as
$q_{\alpha}(X)=\inf\{x\in\mathcal{X}:\mathbb{P}(X\leq x)\geq\alpha\}.$
We denote the distribution function of $X$ as $F$. The $\alpha$-quantile is then the inverse of $F$; that is, $q_{\alpha}(X)=F^{-1}(\alpha)$.
For a set of $n$ elements $Z=\{z_{1},\dots,z_{n}\}$, the left $\alpha$-empirical quantile is given by
$\widehat{Q}^{-}_{\alpha}(Z)=\text{the }\left\lfloor(n+1)\alpha\right\rfloor\text{-th smallest value of }Z,$
and the right $\alpha$-empirical quantile, or the $(1-\alpha)$-empirical quantile, by
$\widehat{Q}^{+}_{\alpha}(Z)=\text{the }\left\lceil(n+1)(1-\alpha)\right\rceil\text{-th smallest value of }Z=-\widehat{Q}^{-}_{\alpha}(-Z),$
where $-Z=\{-z_{1},\dots,-z_{n}\}$. Note that the $n+1$ term serves as a finite sample correction.
p-values   A p-variable [18] is a random variable $P^{\prime}$ such that
$\mathbb{P}\{P^{\prime}\leq\alpha\}\leq\alpha\quad\forall\alpha\in(0,1).$ (1)
The values taken by a p-variable are called p-values. In hypothesis testing, p-values represent the probability of observing results at least as extreme as the ones obtained, assuming the null hypothesis is true. When the inequality becomes equality, the p-variable is precise (rather than conservative) and follows a $\mathcal{U}(0,1)$ distribution.
Conformal prediction
Suppose we have training data $\{(X_{1},Y_{1}),\dots(X_{n},Y_{n})\}$ and a test point $(X_{n+1},Y_{n+1})$ with unknown label $Y_{n+1}$. Let $\mathcal{P}$ denote the joint distribution on $(X,Y)\in\mathcal{X}\times\mathcal{Y}$. Then, assuming that the training and test data are exchangeable, or that the distribution of the training and test data is permutation-invariant, conformal prediction can be used to construct a valid prediction set for $X_{n+1}$ with no further assumptions on $\mathcal{P}$.
Split conformal prediction, a popular variant, starts by splitting the training data indices into disjoint “proper” training and calibration subsets so that $\{1,\dots,n\}=S_{\text{train}}\cup S_{\text{cal}}$. We fit a predictor to the subset $S_{\text{train}}$, so that $\widehat{\mu}=\mathcal{A}(\{(X_{i},Y_{i})\}_{i\in S_{\text{train}}})$, where $\mathcal{A}$ is a model fitting algorithm. Then, given a nonconformity score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, we can compute the scores $R_{i}=s(X_{i},Y_{i})$, or the values of the score function that characterize how nonconformal a label $Y_{i}$ is from its predicted value $\widehat{\mu}(X_{i})$. We compute the scores on the subset $S_{\text{cal}}$. At a given significance level $\alpha$, this procedure allows us to define the prediction set
$\widehat{C}_{\alpha}^{\text{split}}(X_{n+1})=\left\{y\in\mathcal{Y}:s(X_{n+1},y)\leq\widehat{Q}^{+}_{\alpha}(\{s(X_{i},Y_{i})\}_{i\in S_{\text{cal}}})\right\}.$
By construction, the split conformal prediction set results in the (marginal) coverage guarantee
$\mathbb{P}\{Y_{n+1}\in\widehat{C}^{\text{split}}_{\alpha}(X_{n+1})\}\geq 1-\alpha\quad\forall\alpha\in(0,1).$
In other words, the prediction set $\widehat{C}^{\text{split}}_{\alpha}(X_{n+1})$ is valid, or guaranteed to contain the true label $Y_{n+1}$ with probability at least $1-\alpha$.
As an alternative perspective, the quantile can be replaced with the p-value function
$\widehat{p}(x,y)=\frac{1+\sum_{i\in S_{\text{cal}}}\mathbbm{1}\{s(x,y)<s(X_{i},Y_{i})\}}{|S_{\text{cal}}|+1},$
which returns the proportion of calibration points that are less conformal than some test point $(x,y)$, with a finite sample correction. (Note that for an infinite number of samples, this quantity is uniform under exchangeability.) The equivalent prediction set constructed from the p-value function is
$\widehat{C}_{\alpha}^{\text{split}}(X_{n+1})=\left\{y\in\mathcal{Y}:\widehat{p}(X_{n+1},y)>\alpha\right\}.$
Since we only need to fit a single predictor $\widehat{\mu}$, split conformal is computationally cheap. However, it has a statistical cost: if $|S_{\text{cal}}|$ is small, then the variability/randomness in the process used to generate the prediction set is high.
3 Related work
The prediction set for a test point depends on both the data generating process and the choice of model $\widehat{\mu}$; in split conformal, it also depends on how the data is split. This last dependence we can mitigate by repeatedly performing the split conformal procedure and combining the results. In this section, we categorize existing work on combining conformal results into two main approaches: the first focuses on methods for selecting or combining prediction sets, while the second focuses on methods for combining p-values derived from multiple runs of the conformal procedure.
Combining prediction sets
Combining conformal prediction sets was first introduced by Lei et al. [20], who propose combining $K$ split conformal prediction sets by taking their intersection. If each prediction set is constructed at a significance level of $\alpha/K$, then a union bound results in an $\alpha$-level intersection set. Unfortunately, the individual prediction sets (constructed to cover $1-\alpha/K$ of the data) tend to be inefficient—that is, their width or volume is large. Although taking the intersection of these sets reduces the size of the resulting prediction set, the authors demonstrate that, under general conditions, the intersection is wider than the individual prediction sets with probability tending to 1 with an asymptotic number of samples.
Rather than combine prediction sets, Yang and Kuchibhotla [21] propose to select the predictor, from a set of $K$ predictors, that returns the best prediction set. They present two methods: one to select the predictor that gives the most efficient prediction set, but with only approximate validity, and one to select the predictor with the most valid prediction set (with minimal coverage slack), albeit with a width only close to the minimum. Liang et al. [22] point out that the approximate validity of the first method is due to selection bias, and introduce an alternative approach that both selects the most efficient predictor and uses it to construct a valid prediction set.
A recent line of work explores combining conformal prediction sets by majority vote with a $1-2\alpha$ coverage guarantee [23, 24, 17]. In particular, Gasparin and Ramdas [17] characterize the width of the majority vote set and introduce many extensions, including a weighted majority vote method that incorporates prior information in the weights. Our work on combining conformal results by weighted p-values is closely related to this method. However, our formulation allows us to leverage the results of Vovk and Wang [18] to improve the coverage guarantee beyond the standard $1-2\alpha$.
Combining p-values
A substantial body of research has been devoted to developing methods for combining p-values. Here, we focus specifically on approaches that accommodate arbitrary dependence, with a particular emphasis on their applications to conformal prediction. (For a more comprehensive review, see Balasubramanian et al. [25] and DiCiccio et al. [26].) An early example of combining p-values with arbitrary dependence is the Bonferroni method, where the minimum of a set of p-values is scaled by the number of p-values. This method has multiple extensions [27, 28], and was first applied to conformal prediction by Lei et al. [20]. Rüschendorf [29] finds that twice the average of p-values is a p-value; this was later extended to a more general notion of average by Vovk and Wang [18]. Stutz et al. [30] use the result of Rüschendorf [29] to get a $1-2\alpha$ guarantee for the average of p-values, but also propose a novel transformation to get a p-value average with improved coverage. Liu and Xie [31] propose the Cauchy combination test, where the test statistic is a weighted sum of Cauchy-transformed p-values. Wu et al. [32] apply the Cauchy combination test to conformal prediction, and show that their method yields asymptotically exact prediction sets under a weak stability assumption and a joint distribution assumption.
4 Combining conformal prediction sets by weighted p-value
In this section, we extend the method of combining p-values, as proposed by Vovk and Wang [18], to the setting of conformal prediction. Suppose we have $K$ predictors, denoted by $\widehat{\mu}_{1},\dots,\widehat{\mu}_{K}$, where each predictor $\widehat{\mu}_{k}$ is associated with a nonconformity score function $s_{k}$ and calibration set $S_{k}$. For each predictor, we define a p-value function $\widehat{p}_{k}(x,y)$ that measures how well a candidate label $y$ conforms to the predicted outcome for a given $x$. Specifically, for each $k$, the p-value for a given point is given by
$\widehat{p}_{k}(x,y)=\frac{1+\sum_{i\in S_{k}}\mathbbm{1}\{s_{k}(x,y)<s_{k}(X_{i},Y_{i})\}}{|S_{k}|+1}.$ (2)
The prediction set for $X_{n+1}$ is then
$\widehat{C}_{\alpha,k}^{\text{split}}(X_{n+1})=\{y\in\mathcal{Y}:\widehat{p}_{k}(X_{n+1},y)>\alpha\},$
with a guaranteed marginal coverage
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{\alpha,k}^{\text{split}}(X_{n+1})\right\}\geq 1-\alpha\quad\forall\alpha\in(0,1).$ (3)
Now, consider assigning a weight $v_{k}$ to each p-value function $\widehat{p}_{k}$. The weighted average p-value function is then
$\bar{p}(x,y)=\sum_{k=1}^{K}v_{k}\widehat{p}_{k}(x,y)\quad\text{where }\sum_{k=1}^{K}v_{k}=1,$ (4)
with prediction set
$\widehat{C}_{\alpha}^{\text{avg}}(X_{n+1})=\{y\in\mathcal{Y}:\bar{p}(X_{n+1},y)>\alpha\}.$ (5)
We now provide a coverage guarantee for this aggregated prediction set.
Proposition 4.1.
Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$, where $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Then, the prediction set $\widehat{C}_{\alpha}^{\text{avg}}(X_{n+1})$ (5) defined by the weighted average p-value function (4) has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\min\left\{\frac{1}{v},2\right\}\alpha\quad\forall\alpha\in(0,1),$ (6)
where $v=\max\{v_{1},v_{2},\dots,v_{K}\}$ is the largest weight assigned to any of the p-values.
Weighted aggregation provides more flexible coverage guarantees
In a sense, this result provides a spectrum of coverage guarantees based on the weight distribution among the models. If one predictor dominates (i.e., $v>1/2$), the coverage guarantee improves beyond the standard $1-2\alpha$ of typical aggregation methods, allowing the overall coverage to approach that of the most influential predictor. In the extreme case, when all the weight is assigned to a single model ($v=1$), we recover the standard split conformal guarantee of $1-\alpha$ coverage for that model. Thus, we can interpolate between the standard $1-2\alpha$ guarantee of combined models and the $1-\alpha$ guarantee of individual models, with the weights controlling the trade-off. This improves the coverage guarantee for asymmetric aggregation, and opens up the utility of asymmetric generalizations of the many established symmetric aggregation methods (e.g. set-weighted versions of cross-conformal, CV+, etc.)
Since the work of Vovk and Wang holds for arbitrarily dependent p-values, this method is robust across a wide range of scenarios. In practice, this weighting scheme allows users to incorporate prior knowledge about the relative quality of different predictors [18, 17]. The weighting can reflect, for example, expert insights on which of $K$ models should be prioritized as being more reliable.
From p-values to prediction sets   The work of Vovk and Wang is central to our result, allowing us to generalize and improve upon existing conformal guarantees. Still, despite its broad applicability, their work remains underexplored in conformal literature, and the weight-dependent coverage result has not yet been applied to conformal prediction sets. [^1] To the best of our knowledge, Gasparin and Ramdas [17] is the only work so far to use asymmetric set aggregation, and they derive the standard $1-2\alpha$ guarantee of symmetric aggregation methods. We attribute this oversight to several factors.
Vovk and Wang frame their work in terms of merging functions and p-values, without reference to prediction sets or conformal prediction. As a result, subsequent research has similarly focused on p-values and related test statistics, and has limited connection to conformal literature. Meanwhile, conformal literature typically uses a quantile-based construction of prediction sets, rather than a p-value construction, and work in conformal prediction set aggregation typically operates directly on the sets, rather than working with the associated p-value functions. These trends may contribute to the pattern where p-value results are not always propagated to the wider conformal prediction community. Our result shows that the p-value presentation offers unique benefits to prediction set aggregation, and we hope that our work may encourage renewed interest in the connection between p-values and conformal prediction.
5 Extending to data-dependent weights
In many scenarios, it is desirable to assign model weights based on the observed data. Data-dependent weights offer an adaptive approach that adjusts the influence of each model according to specific characteristics of the data. However, when weights depend on the data, they also become dependent on the p-variables; as a result, the theory of Vovk and Wang [18] does not directly apply to this setting. Nevertheless, we can build on their approach to propose a simple, empirical method that allows us to recover a coverage guarantee, even when the weights depend on the data.
The key idea is to directly use the definition of a p-variable (1). The sum of weighted p-variables, where the weights are dependent on the p-variables, is not necessarily a p-variable. However, the weighted sum can become a p-variable by a linear transformation that both satisfies the definition of a p-variable and preserves the proportions of the weights.
Let the p-variables of the predictors be $P_{1},\dots,P_{K}\in\mathcal{U}$, where $\mathcal{U}$ is the set of all uniformly distributed random variables. The weights are given by a random vector $\mathbf{W}=(W_{1},\dots,W_{K})$ in the $(K-1)$-dimensional simplex $\Delta_{K-1}:=\{\mathbf{w}=(w_{1},\dots,w_{K})\in[0,1]^{K}:w_{1}+\dots+w_{K}=1\}$ depending on the data $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$. We define the weighted average function
$p_{\text{all}}(x,y;\mathbf{w})=\sum_{k}w_{k}\widehat{p}_{k}(x,y),$ (7)
giving associated random variable $P_{\text{all}}:=p_{\text{all}}(X,Y;\mathbf{W})$ with distribution $F_{P_{\text{all}}}$. Then, we propose to learn the scalar
$m^{*}$ $=\inf\{m\in\mathbb{R}^{+}:\mathbb{P}\{mP_{\text{all}}\leq\alpha\}\leq\alpha\quad\forall\alpha\in(0,1)\}$ (8)
$=\inf\{m\in\mathbb{R}^{+}:F_{P_{\text{all}}}(\alpha/m)\leq\alpha\quad\forall\alpha\in(0,1)\}=\sup_{\delta\in(0,\infty)}\frac{F_{P_{\text{all}}}(\delta)}{\delta}.$
(See §A.4 for a full derivation.) The scaling defined in (8) transforms the weighted average $P_{\text{all}}$ into a valid p-variable, recovering a coverage guarantee. We formalize this in the following result.
Proposition 5.1.
Let $\widehat{C}_{\alpha,1}^{\text{split}}(X_{n+1}),\dots,\widehat{C}_{\alpha,K}^{\text{split}}(X_{n+1})$ be $K$ prediction sets defined by p-value functions $\widehat{p}_{1},\dots,\widehat{p}_{K}$ (2) on $X_{n+1}$ corresponding to p-variables $P_{1},\dots,P_{K}$. Suppose $1-\alpha$ coverage (3) holds for each set $k\in[K]$. Let $(W_{1},\dots,W_{K})$ be a random vector in $\Delta_{K-1}$ depending on $\{(X_{i},Y_{i})\}_{i\in[n]}\cup\{X_{n+1}\}$. Then, the prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ defined by the weighted average $p_{\text{all}}$ (7) corresponding to random variable $P_{\text{all}}$ has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-m^{*}\alpha\quad\forall\alpha\in(0,1),$ (9)
where $m^{*}$ is defined in (8).
Algorithm 1 Constructing a weighted aggregate of prediction sets with valid coverage
1:Input: $K$ predictors from potentially overlapping datasets, $K$ weights dependent on the predictors (i.e. potentially learned from the same data), data with indices $\{1,\dots,n\}=S_{\text{cal}}\cup S_{\text{merge}}$, and test example $X_{n+1}$.
2:Output: Prediction set around the test example $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ with coverage $1-\widehat{m}^{*}\alpha$.
3:
4:Step 1. Derive the p-value function $p_{\text{all}}$.
5: 1.1. Using calibration set $S_{\text{cal}}$, derive p-value functions $\widehat{p}_{k}$ for each of the predictors (2).
6: 1.2. Take the weighted sum of the p-value functions $\widehat{p}_{k}$ with the $K$ dependent weights. The result is the function $p_{\text{all}}$, with prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$.
7:
8:Step 2. Compute the correction factor $\widehat{m}^{*}$.
9: 2.1. Using the points of merging set $S_{\text{merge}}$ with the function $p_{\text{all}}$, get samples of $\widehat{F}_{P_{\text{all}}}$.
10: 2.2. Derive the correction factor $\widehat{m}^{*}$ from the samples of $\widehat{F}_{P_{\text{all}}}$ (11).
Computing $m^{*}$ with a merging set
The exact value of $m^{*}$ depends on $F_{P_{\text{all}}}$, which is generally unknown. We therefore propose to set aside a subset of the data $S_{\text{merge}}$ [^2] So named because $S_{\text{merge}}$ is used to learn the correction that makes the merging function [18] for the weighted average empirically precise. to estimate $F_{P_{\text{all}}}$ in practice.
Typically, an $n$-sample empirical CDF for $P_{\text{all}}$ would be given by
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in\mathcal{S}_{\text{merge}}}\mathbbm{1}\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\}}{|\mathcal{S}_{\text{merge}}|};$
we use the more conservative function
[…]
as it tends to yield more stable estimates in finite samples. The correction factor is then computed as
$\widehat{m}^{*}=\max_{i\in\mathcal{S}_{\text{merge}}}\frac{\widehat{F}^{\text{cons}}_{P_{\text{all}}}(p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)}))}{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})},$ (11)
following (8). (For more detail on why (11) is equivalent to (8) for an empirical CDF, see §A.5.) The computation of $\widehat{m}^{*}$ yields the following finite-sample coverage guarantee:
Proposition 5.2.
(Finite-sample guarantee.) Under the same assumptions as Proposition 5.1, fix a user-chosen failure probability $\delta\in(0,1)$ and set
$\varepsilon=\sqrt{\frac{\log(2/\delta)}{2|S_{\text{merge}}|}}.$
The prediction set $\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})$ defined by the weighted average $p_{\text{all}}$ (7) has the following coverage guarantee:
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\right)\quad\forall\alpha\in(0,1),$ (12)
where $\widehat{m}^{*}$ is defined in (11).
In §B.1.1, we explore how the size of $S_{\text{merge}}$ affects $\widehat{F}^{\text{cons}}_{P_{\text{all}}}$ and therefore, the quality of the correction $\widehat{m}^{*}$ and the tightness of the coverage guarantee. We find that in practice, a modest number of samples ($<200$) is sufficient to achieve accurate coverage. We summarize our procedure in Algorithm 1.
Data-dependent weights give a form of conditional coverage
Data-dependent weights allow the influence of each model to be adjusted based on how well it performs for a specific data point. This approach enables the construction of aggregated prediction sets that are tailored to the characteristics of the given data, which can be viewed as a form of locally conditional coverage. To be clear, true X-conditional coverage—where the coverage guarantee is conditioned on the current input—is impossible without additional distributional assumptions [33, 34]. However, we demonstrate in §6 that data-dependent weights allow us to create aggregated prediction sets with data-adaptive coverage that can greatly improve conditional validity in practice.
Achieving tighter guarantees
Our scaling correction factor $m^{*}$ provides a coverage guarantee that holds for all significance levels $\alpha\in(0,1)$. However, guarantees on coverage for all $\alpha$ can lead to overly conservative prediction sets, which may be unnecessarily restrictive in practice when we do not need guarantees for every possible significance level. Thus, we propose two alternatives: for a specific significance level $\alpha^{\prime}$, we can learn a correction factor
$m^{\dagger}=\inf\{m\in\mathbb{R}^{+}:\mathbb{P}\{mP_{\text{all}}\leq\alpha\}\leq\alpha\quad\forall\alpha\in(0,\alpha^{\prime}]\},$ (13)
or the even stricter
$m^{\ddagger}=\inf\{m\in\mathbb{R}^{+}:\mathbb{P}\{mP_{\text{all}}\leq\alpha^{\prime}\}\leq\alpha^{\prime}\}.$ (14)
Note that (13) is not without precedent: for the $1-2\alpha$ coverage guarantees of aggregation methods like cross-conformal and jackknife+, $\alpha^{\prime}=0.5$ is the highest significance level of interest.
6 Application: Mixture-of-Experts
Mixture-of-Experts (MoE) is a machine learning framework designed to combine the predictions of multiple specialized models, called experts [19]. Each expert in an MoE focuses on unique aspects of the problem by learning different representations of the input data. A central component of this framework is the routing network, which determines how to combine the experts’ outputs. Specifically, for an input $x$, the routing network of a traditional (soft) MoE assigns weights $W_{k}(x)$ to each expert output $f_{k}(x)$, producing the final prediction as a weighted sum
$f(x)=\sum_{k}W_{k}(x)f_{k}(x).$ (15)
By learning how to route different inputs to the most appropriate experts, the routing network implicitly conditions the model’s final prediction on the combination of experts that fits the given input. In this way, the routing network can be viewed as learning a form of conditional relationship between the input features and the expertise of each model.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 1: Left: Network diagram for MoE. For traditional MoE, the aggregation module takes a weighted sum of the outputs from each expert. To learn weight-dependent prediction sets, we instead propose to combine the prediction sets of each expert by weighted p-value. Right, top row: Comparison of split conformal prediction sets with those learned from weighted aggregation. Weighted aggregation allows overall coverage to follow the coverage of the dominant expert, rather than remain purely marginal. Right, bottom row: Another comparison of split conformal with weighted aggregation, with the latter showing local coverage with smooth transitions.
The routing network’s ability to learn data-dependent weights for each expert makes the MoE framework a natural setting for applying our method of aggregating prediction sets by weighted p-values. We collect the routing weights into the simplex-valued vector $W(x):=(W_{1}(x),\dots,W_{K}(x))$; then, for a new input $X_{n+1}$, the routing network outputs the weight vector $\mathbf{W}^{(n+1)}:=W(X_{n+1})$, whose $k$th component we write as $W^{(n+1)}_{k}$. Denote the p-value function of the $k$th expert as $\widehat{p}_{k}$. The weighted aggregate p-value function for some label $y$ is then
$p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})=\sum_{k}W_{k}^{(n+1)}\widehat{p}_{k}(X_{n+1},y).$
At a significance level $\alpha$, the associated prediction set is
$\widehat{C}_{\alpha}^{\text{data-avg}}(X_{n+1})=\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\}.$
Intuitively, the weights assigned by the routing network adjust the p-value contribution of each expert by how predictive they are for a given point. Data-dependent weights enable data-adaptive coverage, which is more locally robust than split conformal with the full MoE (15) as a predictor (Figure 1).
Baselines
We refer to our proposed method of aggregating expert p-value functions with learned weights as weighted aggregation. As a baseline, we compare against split conformal with the full MoE predictor, and we refer to this simply as split conformal for brevity.
We compare against conformal quantile regression (CQR) [35] for experiments evaluating the local validity of our method. Weighted aggregation can be applied on top of conformal quantile regression, so we also evaluate a hybrid method combining the two. To assess performance, we consider marginal coverage, worst-slice (WS) coverage [36], and prediction set size.
We consider two practical variants of weighted aggregation:
  • WA targeted $(0,\alpha^{\prime}]$: Coverage is guaranteed for all $\alpha\in(0,\alpha^{\prime})$ using the $m^{\dagger}$ correction (13).
  • WA precise $\alpha^{\prime}$: Coverage is guaranteed for $\alpha=\alpha^{\prime}$ only, using the $m^{\ddagger}$ correction (14).
Experiment overview
We present our main experimental results on the local validity of our method on real-world data in Figure 2. These results focus on regression tasks, where we can compare the standard absolute residual and CQR nonconformity scores; we evaluate our method on classification tasks in §B.2.1. To complement this analysis, which focuses on WS coverage, we also assess local validity from another perspective, by examining coverage disparities across demographic groups in the Communities and Crimes dataset [37] in §B.2.2.
In the Appendix, we analyze how various factors impact coverage and prediction set size in a series of ablative studies on synthetic data. We investigate: the effects of merging set size $|S_{\text{merge}}|$B.1.1), shared feature information (§B.1.2), and the different weighted aggregation variants (§B.1.3).
The following experiments are conducted with an MoE where the routing network and experts are all linear models for simplicity. All results are averaged over 200 trials.
Results
Figure 2: Local validity experiments comparing split conformal to weighted aggregation (WA) using absolute residual scores and CQR scores. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, and prediction set size from left to right. WA consistently improves WS coverage across datasets, which split conformal undercovers.
In Figure 2, we compare split conformal to weighted aggregation using both absolute residual scores and CQR scores. From the marginal coverage plots (left column), we see that split conformal consistently achieves coverage closest to the nominal $1-\alpha$ level, while the weighted aggregation methods tend to overcover. However, the WS coverage plots (middle column) highlight a well-known limitation of split conformal: it does not maintain coverage uniformly across the data, significantly undercovering on the WS slab. While CQR scores improve split conformal WS coverage on some datasets (Airfoil and Communities), they are still insufficient to close the coverage gap.
Weighted aggregation significantly improves WS coverage, coming close to or meeting the nominal level on all datasets. In addition, the gap between marginal and WS coverage is notably smaller for weighted aggregation than for split conformal, indicating that weighted aggregation provides more uniform coverage, including in challenging regions of the data. These results suggest that weighted aggregation is useful in applications where local validity is a priority.
The prediction set size plots (right column) illustrate the standard trade-off between coverage and efficiency, where higher-coverage methods like WA targeted tend to produce larger prediction sets than lower-coverage methods like split conformal. According to these results, we suggest WA targeted if maintaining coverage in challenging regions is the primary concern. If both efficiency and local validity are important, WA precise provides a reasonable middle ground—offering improved WS coverage over split conformal, while also maintaining more compact prediction set size.
7 Conclusion
The asymmetric (weighted) aggregation of prediction sets is a flexible generalization of the more standard symmetric aggregation. Intuitively, coverage for this case should not be constant, but should depend on the distribution of weights. The results of Vovk and Wang [18] enable us to achieve this for conformal prediction, so that for prior weights using expert information, we improve the coverage guarantee when the distribution of weights is sufficiently asymmetric. In addition, we propose a method for data-dependent weights (e.g. weights learned from data) where coverage depends on the weight distribution. We find that allowing for data-dependent weights provides a form of conditional coverage, which we verify in experiments on WS coverage and coverage across demographic groups.
Acknowledgements
GW and RC are partially supported by a ONR MURI grant N00014-20-1-2781. DP is partially supported by an Amazon AI2AI fellowship. DP, SS, and AL are partially supported by the Gordon and Betty Moore Foundation grant #12128. AL is also supported by an Amazon Research Award.
Appendix A Proofs and additional theoretical details
A.1 Proof of Proposition 4.1
Following the notation of Vovk and Wang [18], we define the merging function
$M_{1,\mathbf{v}}(p_{1},\dots,p_{K})=(v_{1}p_{1}+\dots+v_{K}p_{K})\quad\text{where }\sum_{k=1}^{K}v_{k}=1.$
Recall that we use p-variables to represent our p-value functions applied to data; that is, $P_{k}=\widehat{p}_{k}(X,Y)$. Then $\bar{P}=M_{1,\mathbf{v}}(P_{1},\dots,P_{K})$. By Proposition 9 of Vovk and Wang [18], $A_{\mathbf{v}}M_{1,\mathbf{v}}$ is a precise merging function, where $A_{\mathbf{v}}=\min\left\{\frac{1}{v},2\right\}$ and $v=\max\{v_{1},...,v_{K}\}$. Thus, $A_{\mathbf{v}}\bar{P}$ is a p-variable, and $\mathbb{P}\left\{A_{\mathbf{v}}\bar{P}\leq\alpha\right\}\leq\alpha$, or
$\mathbb{P}\left\{\bar{P}\leq\alpha\right\}\leq A_{\mathbf{v}}\alpha=\min\left\{\frac{1}{v},2\right\}\alpha.$
Under exchangeability, the prediction set
$\widehat{C}_{\alpha}^{\text{avg}}(X_{n+1})=\{y\in\mathcal{Y}:\bar{p}(X_{n+1},y)>\alpha\}$
constructed from $\bar{p}$ has the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\min\left\{\frac{1}{v},2\right\}\alpha.$
A.2 Proof of Proposition 5.1
Proposition 2 follows from the definition of the correction factor $m^{*}$ (8).
We define $m^{*}$ to be the smallest positive value such that $m^{*}P_{\text{all}}$ is a p-variable, meaning that
$\mathbb{P}\left\{m^{*}P_{\text{all}}\leq\alpha\right\}\leq\alpha.$
We can rearrange this condition to be
$\mathbb{P}\left\{P_{\text{all}}\leq\frac{\alpha}{m^{*}}\right\}\leq\alpha;$
if we define $\alpha^{\prime}=\alpha/m^{*}$, this becomes
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\right\}\leq m^{*}\alpha^{\prime},$
or, equivalently,
$\mathbb{P}\left\{P_{\text{all}}>\alpha^{\prime}\right\}\geq 1-m^{*}\alpha^{\prime}.$
Recall that $P_{\text{all}}:=p_{\text{all}}(X,Y,\mathbf{W})$. Then under exchangeability, the prediction set
$\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
satisfies the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-m^{*}\alpha.$
Note that other transformations of $P_{\text{all}}$ can also yield valid p-variables while preserving the relative weighting. For example, if $P_{\text{all}}$ is shifted rather than scaled, then the condition
$\mathbb{P}\left\{m^{\prime}+P_{\text{all}}\leq\alpha\right\}\leq\alpha$
leads to a coverage guarantee of $1-(m^{\prime}+\alpha)$. We focus on the scale correction factor $m^{*}$ because it aligns with the framework of Vovk and Wang [18] and works well in practice.
A.3 Proof of Proposition 5.2
Let $\mathcal{G}$ be the event that the worst-case distance between the true CDF $F_{P_{\text{all}}}$ and the empirical CDF $\widehat{F}_{P_{\text{all}}}$ is bounded by some maximum allowable deviation; that is,
$\mathcal{G}=\left\{\sup_{x\in\mathbb{R}}\left|\widehat{F}_{P_{\text{all}}}(x)-F_{P_{\text{all}}}(x)\right|\leq\varepsilon\right\}\quad\text{where }\varepsilon=\sqrt{\frac{\log(2/\delta)}{2|S_{\text{merge}}|}}$
for some user-chosen failure probability $\delta\in(0,1)$. By the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality,
$\mathbb{P}(\mathcal{G})\geq 1-\delta.$
Now, suppose we have a fixed merging set $S_{\text{merge}}$ such that $\mathcal{G}$ holds. By definition of a CDF,
$\mathbb{P}\left\{\widehat{m}^{*}P_{\text{all}}\leq\alpha\mid S_{\text{merge}}\right\}=F_{P_{\text{all}}}\left(\frac{\alpha}{\widehat{m}^{*}}\right),$ (16)
and under the same fixed $S_{\text{merge}}$,
$F_{P_{\text{all}}}(x)\leq\widehat{F}_{P_{\text{all}}}(x)+\varepsilon.$
We take $x=\alpha/\widehat{m}^{*}$; this allows us to express the bound as
$F_{P_{\text{all}}}\left(\frac{\alpha}{\widehat{m}^{*}}\right)\leq\widehat{F}_{P_{\text{all}}}\left(\frac{\alpha}{\widehat{m}^{*}}\right)+\varepsilon.$ (17)
By construction of $\widehat{m}^{*}$ (11), we know that $\widehat{F}_{P_{\text{all}}}(x)/x\leq\widehat{m}^{*}$ for all $x$. Substituting again $x=\alpha/\widehat{m}^{*}$, we have
$\widehat{F}_{P_{\text{all}}}\left(\frac{\alpha}{\widehat{m}^{*}}\right)\leq\widehat{m}^{*}\frac{\alpha}{\widehat{m}^{*}}=\alpha.$ (18)
Combining (17) and (18) gives
$F_{P_{\text{all}}}\left(\frac{\alpha}{\widehat{m}^{*}}\right)\leq\alpha+\varepsilon.$
Applying this to (16) gives
$\mathbb{P}\left\{\widehat{m}^{*}P_{\text{all}}\leq\alpha\mid S_{\text{merge}}\right\}\leq\alpha+\varepsilon,$
or
$\mathbb{P}\left\{P_{\text{all}}\leq\alpha^{\prime}\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha^{\prime}+\varepsilon$
by a change of variable.
Recall that $P_{\text{all}}:=p_{\text{all}}(X,Y,\mathbf{W})$. Under exchangeability, the prediction set
$\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})=\left\{y\in\mathcal{Y}:p_{\text{all}}(X_{n+1},y;\mathbf{W}^{(n+1)})>\alpha\right\}$
has a conditional miscoverage guarantee of
$\mathbb{P}\left\{Y_{n+1}\notin\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\mid S_{\text{merge}}\right\}\leq\widehat{m}^{*}\alpha+\varepsilon.$
Let us denote the miscoverage indicator as
$\text{MC}=\mathbbm{1}\left\{Y_{n+1}\notin\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}.$
The indicator random variable allows us to easily switch between probability and expectation in order to marginalize over $S_{\text{merge}}$:
$\mathbb{P}\left\{\text{MC}=1\right\}$ $=\mathbb{E}\left[\text{MC}\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\text{MC}\mid S_{\text{merge}}\right]\right]$
$=\mathbb{E}\left[\mathbb{P}\left\{\text{MC}=1\mid S_{\text{merge}}\right\}\right]$
$\leq\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon\right)\mathbb{P}(\mathcal{G})+1\cdot\mathbb{P}(\mathcal{G}^{c})$
$\leq\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta.$
This gives us the coverage guarantee
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}^{\text{data-avg}}_{\alpha}(X_{n+1})\right\}\geq 1-\left(\alpha\mathbb{E}[\widehat{m}^{*}]+\varepsilon+\delta\right).$
Note that the same argument applies directly if we replace the empirical CDF $\widehat{F}_{P_{\text{all}}}$ by the conservative version $\widehat{F}_{P_{\text{all}}}^{\text{cons}}$ (10) because the two differ by at most $(|\mathcal{S}_{\text{merge}}|+1)^{-1}$, a deterministic offset that can simply be absorbed into the DKW tolerance by replacing $\varepsilon$ with $\varepsilon+(|\mathcal{S}_{\text{merge}}|+1)^{-1}$.
A.4 More detailed derivation of $m^{*}$
For some distribution function $F_{P_{\text{all}}}$, we define
$m^{*}=\inf\left\{m>0:F_{P_{\text{all}}}(\alpha/m)\leq\alpha\text{ for all }\alpha\in(0,1)\right\},\qquad c=\sup_{\delta>0}\frac{F_{P_{\text{all}}}(\delta)}{\delta}.$
We aim to prove that $m^{*}=c$ to establish the equivalence in (8). To aid our proof, we define the feasible set
$S=\left\{m>0:F_{P_{\text{all}}}(\alpha/m)\leq\alpha\text{ for all }\alpha\in(0,1)\right\},$
where $m^{*}=\inf(S)$.
We begin by showing that $m^{*}\leq c$. To this end, consider any $m>c$. By definition of $c$,
$F_{P_{\text{all}}}(\delta)\leq c\,\delta<m\,\delta\quad\forall\delta>0.$
Pick $\delta=\alpha/m$ for any $\alpha\in(0,1)$ so that $\delta\in(0,1/m)$. Substituting yields
$F_{P_{\text{all}}}(\alpha/m)<m(\alpha/m)=\alpha.$
This shows that every $m>c$ satisfies the feasibility condition, so every $m>c$ is in $S$; it follows that $S\supset(c,\infty)$. Then, as the infimum of $S$, $m^{*}\leq c$.
To show that $m^{*}\geq c$, consider any feasible $m>0$, i.e. assume
$F_{P_{\text{all}}}(\alpha/m)\leq\alpha\quad\forall\alpha\in(0,1).$
For any $\delta>0$, there are two cases.
  1. 1.
    If $\delta\geq 1/m$, since $F_{P_{\text{all}}}(\delta)\leq 1$,
    $F_{P_{\text{all}}}(\delta)/\delta\leq 1/\delta\leq m.$
  2. 2.
    If $\delta\in(0,1/m)$, let us select $\alpha=m\delta\in(0,1)$. By feasibility of $m$,
    $F_{P_{\text{all}}}(\delta)=F_{P_{\text{all}}}(\alpha/m)\leq\alpha=m\delta,$
    or
    $F_{P_{\text{all}}}(\delta)/\delta\leq m.$
The two cases above establish $m\geq c$ for every $m\in S$, or that $S\subset(c,\infty)$. Thus, $c$ is a lower bound of $S$ and $m^{*}\geq c$.
A.5 Computing $\widehat{m}^{*}$ from an empirical CDF
For random variable $P_{\text{all}}=p_{\text{all}}(X,Y;\mathbf{W})$, a merging correction factor $\lambda$ is a positive scalar that ensures that a $\lambda P_{\text{all}}$ is a valid p-value. That is,
$\lambda\in\{m>0:\mathbb{P}\{mP_{\text{all}}\leq\alpha\}\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}=\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}.$
Scaling by a merging correction factor gives us the guarantee
$\mathbb{P}\{P_{\text{all}}>\alpha\}=1-\lambda\alpha.$
To achieve the tightest guarantee (with application to all $\alpha\in(0,1)$), we define the minimal merging correction factor $m^{*}$ to be
$m^{*}=\inf\{m>0:F_{P_{\text{all}}}\left(\alpha/m\right)\leq\alpha\leavevmode\nobreak\ \forall\alpha\in(0,1)\}=\sup_{\delta>0}{\frac{F_{P_{\text{all}}}(\delta)}{\delta}}.$ (19)
(Equation (19) is a restatement of (8) from the main body; we replicate it here for easy reference.)
In practice, we compute the minimal merging factor $\widehat{m}^{*}$ by first constructing the empirical CDF of $P_{\text{all}}$ over the merging set
$\widehat{F}_{P_{\text{all}}}(\alpha)=\frac{\sum_{i\in S_{\text{merge}}}\mathbbm{1}\left\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\leq\alpha\right\}}{|S_{\text{merge}}|},$
or its conservative version $\widehat{F}^{\text{cons}}_{P_{\text{all}}}$ (10). For the sake of conciseness, let
$\widehat{F}_{i}:=\widehat{F}_{P_{\text{all}}}\left(p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})\right).$
Since $\widehat{F}_{P_{\text{all}}}$ is a right-continuous step function that only jumps at the observed values $\{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)}):i\in S_{\text{merge}}\}$, the supremum in (19) is attained at one of these points. In fact, for any $\delta$ not equal to one of these values, shifting $\delta$ slightly to the right (toward the next jump) increases the denominator without changing the numerator, decreasing the ratio. For this reason, it suffices to take the maximum over the observed p-values, yielding the empirical merging factor
$\widehat{m}^{*}=\max_{\begin{subarray}{c}i\in S_{\text{merge}}\\[1.0pt] p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})>0\end{subarray}}\frac{\widehat{F}_{i}}{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})}.$
In some cases, we are only interested in coverage above a certain significance level $\alpha$ as in (13). Then we only need to find the first point on the empirical CDF that surpasses $\alpha$, or
$\bar{\alpha}=\operatornamewithlimits{min}_{\begin{subarray}{c}i\in S_{\text{merge}}\\ \widehat{F}_{i}\geq\alpha\end{subarray}}\widehat{F}_{i},$
and then compute the scaling factor to be
$\widehat{m}^{\dagger}=\max_{\begin{subarray}{c}i\in S_{\text{merge}}\\[1.0pt] p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})>0\\ \widehat{F}_{i}\leq\bar{\alpha}\end{subarray}}\frac{\widehat{F}_{i}}{p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})}.$
Appendix B Additional experiments
B.1 Synthetic data
We generate a simple homoskedastic dataset to simulate a regression task, where each input is a 16-dimensional Gaussian with random parameters, and the label is the sum of the different dimensions with additive noise (see §E.1 for more detail). We use this dataset to investigate how various factors impact the coverage and interval width of our prediction sets.
B.1.1 Coverage and size of prediction sets improve with larger $|S_{\text{merge}}|$
The merging set $S_{\text{merge}}$ allows us to construct an empirical CDF for $P_{\text{all}}$. To improve stability in finite samples, we apply a conservative correction to the typical formula for the empirical CDF, and use this conservative CDF (10) to compute the correction factor $\widehat{m}^{*}$ (11). Our conservative CDF tends to be overly conservative when $|S_{\text{merge}}|$ is small, leading to overcoverage. However, as $|S_{\text{merge}}|$ grows and the empirical CDF approaches the true CDF, coverage becomes closer to nominal and the prediction sets become less conservative.
Figure 3 illustrates how the size of the merging set $|S_{\text{merge}}|$ impacts mean coverage. On the left plot, we use the $m^{\dagger}$ correction (13) for WA targeted (that is, $(0,\alpha)$ aggregation), and we compare different ways of assigning features to experts. On the right plot, we use a non-overlapping feature assignment, and we compare the different variants of weighted aggregation. Our results show that larger $|S_{\text{merge}}|$ leads to tighter coverage, with 160 samples being sufficient to achieve $<3\%$ overcoverage for most feature assignment methods. Interestingly, configurations where experts have no overlapping features tend to overcover the most. We explore the effect of feature assignment in the next section.
Figure 3: Mean coverage compared to the size of the merging set $S_{\text{merge}}$, with different feature assignments (left) and different weighted aggregation variants (right). Overall, we find that coverage improves as the merging set gets larger for all methods. We note that the merging set does not have to be prohibitively large to produce decent results: for example, most feature assignment methods overcover by only 2% with fewer than 200 samples.
B.1.2 Feature information overlap leads to higher coverage and more efficient prediction sets
To better understand how the allocation of features to experts affects the behavior of MoE weighted aggregation, we define four feature assignment methods (for our MoE of four experts):
  • Features 15/16: each expert predicts from 15 of the 16 available features.
  • Features 12/16: each expert predicts from 12 of the 16 available features.
  • Share 1/2: all experts share 8 of the 16 features and partition the remaining 8 (2 features each).
  • No Overlap: the experts partition the 16 features (4 features each).
Figure 4 shows how different feature assignment methods affect coverage and prediction set size. Broadly, we observe that greater feature overlap leads to higher coverage (exceeding the nominal level) and more efficient prediction sets. In the MoE setting, this may be because feature sharing leads to more consistent estimations across experts, improving the reliability of the aggregated p-values and, in turn, reducing the size of the prediction sets. More generally, this may imply that the information redundancy introduced by feature overlap allows for better sample efficiency. This parallels findings in aggregation methods like cross-conformal and jackknife+, which also tend to produce smaller prediction sets than split conformal by reusing data. These results suggest that feature sharing is an important design consideration when aggregating prediction sets from multiple models.
Figure 4: Coverage (left) and prediction set size (right) for different feature assignment methods and weighted aggregation methods. Our results indicate that sharing fewer features leads to tighter coverage, but sharing more features leads to more efficient (smaller) prediction sets. We also see that WA precise tends to have coverage that is closest to nominal and the most efficient prediction sets, albeit with a much looser guarantee.
B.1.3 More general coverage guarantees result in more conservative prediction sets
We now compare the different variants of weighted aggregation described in §6, focusing on how the generality of the guarantee for each variant affects its empirical coverage.
Figure 4 compares the effects of different feature assignment methods and for each weighted aggregation variant (with $\alpha^{\prime}=0.1$), and Figure 5 shows how coverage varies across $\alpha$ for each variant.
Unsurprisingly, WA precise—which provides the narrowest guarantee, targeting a single $\alpha$—achieves coverage closest to the nominal level. In contrast, WA targeted offers more general guarantees over a range of $\alpha$ values, but tends to overcover, reflecting the conservativeness built into the method to accommodate worst-case behavior. We observe the same pattern on UCI data in Figure 2. These results illustrate the trade-off between the conservativeness of a method and the generality of its guarantee, and suggest that more targeted guarantees may be preferable when tighter coverage is important.
Figure 5: Mean coverage across significance levels $\alpha$ for the different weighted aggregation variants. WA precise achieves close to nominal coverage, while WA targeted tends to overcover (but offers more general guarantees).
B.2 Real data
B.2.1 Weighted aggregation improves local validity for classification
Figure 6: Local validity experiments comparing split conformal (green) to weighted aggregation (orange, purple, and pink) on the classification task. Like in Figure 2, each row corresponds to a dataset, with plots for marginal coverage, WS coverage, and prediction set size from left to right. We find that for classification as well as regression, WA improves WS coverage over split conformal, which undercovers in the WS region.
The experiments in Figure 6 mirror those in Figure 2, but for classification instead of regression. Like with the other set of experiments, we compare split conformal to weighted aggregation, and we see that split conformal achieves marginal coverage close to nominal but consistently undercovers on the WS slab. In contrast, weighted aggregation maintains much better WS coverage, suggesting it offers better local validity in the classification setting as well.
As before, we note that there is trade-off between coverage and efficiency. WA targeted is the best choice in terms of WS coverage, but WA precise balances the improved WS coverage of weighted aggregation methods with an efficiency that is closer to the split conformal. However, if marginal coverage and prediction set efficiency are the only priorities, then split conformal should be preferred.
B.2.2 Local validity over demographic groups
Figure 7: Coverage (top) and prediction set size (bottom) for split conformal (green) and weighted aggregation variants (orange, purple, and pink) across subgroups with top 50th percentile racial representation. Split conformal has precise marginal coverage (“All”), but WA variants have more consistent coverage across subgroups, with WA targeted meeting coverage for all subgroups. Error bars represent 95% confidence intervals.
Figure 8: Coverage (top) and prediction set size (bottom) for split conformal and weighted aggregation variants across subgroups with top 70th percentile racial representation. As in Figure 7, WA targeted meets coverage for all subgroups. Error bars represent 95% confidence intervals.
In this section, we evaluate our methods on the Communities and Crimes dataset [37], where the task is to predict the per capita violent crime rate of a community based on its demographic features, and our primary interest is in understanding how coverage differs across communities with varying racial compositions [38].
Figures 7 and 8 compare split conformal with weighted aggregation variants in terms of coverage and prediction set size across demographic groups. Unlike Figures 2 and 6, where local validity is assessed via WS coverage, these experiments evaluate local validity in terms of consistency across demographic groups. Thus, we display group-specific performance for each method, which we group further by the two types of nonconformity score, absolute residual and CQR score.
The demographic groups in Figures 7 and 8 represent communities where a particular racial demographic is in the top $p$-percentile of representation [38]. Figure 7 shows results for $p=50$, and Figure 8 for $p=70$.
Marginal coverage and coverage across groups
The “All” category in the coverage plots represents coverage across all demographic groups, or marginal coverage. As in prior experiments, we see that split conformal achieves marginal coverage closest to nominal, with WA precise close behind, while WA targeted variants tend to overcover. However, although split conformal enjoys precise marginal coverage, it also exhibits substantial disparities in performance across demographic groups, significantly undercovering for Black and Hispanic groups while overcovering for White. In contrast, WA variants display far less demographic variation, with WA targeted achieving coverage for all demographic groups.
Comparing nonconformity scores
For all demographic groups except the Asian group, using CQR scores instead of absolute residuals improves coverage for split conformal, with the improvement being most pronounced for the Black group. However, while CQR reduces undercoverage, it is never sufficient to fully close the coverage gap for an undercovered group. Rather, its primary benefit appears to be in reducing the variability in coverage across groups. Across all methods, the variance in CQR-based coverage is lower than that of absolute residuals, suggesting that CQR contributes to more stable group-wise coverage, even if it does not fully mitigate disparities.
Appendix C Other methods for weighted prediction set aggregation
C.1 Weighted majority vote
The idea to combine conformal prediction sets by weighted majority vote was introduced by Gasparin and Ramdas [17] as an extension of the majority vote method first proposed by Cherubin [23]. At first consideration, weighted majority vote appears to differ from our method: weighted majority vote performs weighted aggregation of prediction sets, while our method performs weighted aggregation of the p-values associated with prediction sets. Despite this distinction, Appendix B of Gasparin and Ramdas [17] shows that these two methods are, in fact, dual to each other. Thus, under data-independent weights, our method is essentially the same as theirs. Nevertheless, our p-value formulation enables two key extensions that go beyond weighted majority vote.
  • Our formulation allows us to apply the result of Vovk and Wang [18] to strengthen coverage guarantees for data-independent weights when they are sufficiently asymmetric.
  • Our formulation provides a principled extension to data-dependent weights by transforming the weighted average of p-variables to also be a valid p-variable. Not only does this allow us to use weights learned from data, but it also yields a form of local validity, a property not available to existing set aggregation methods.
Because our method is a dual formulation to the weighted majority vote method of Gasparin and Ramdas [17], we do not include it as a separate baseline to avoid redundancy.
C.2 Extending the p-variable transformation of Stutz et al. [30] to weighted aggregation
To the best of our knowledge, Gasparin and Ramdas [17] present the only existing method to address weighted prediction set aggregation, and the method of Stutz et al. [30] is designed for the different problem of uncertainty in the ground truth labels. To address their problem, Stutz et al. [30] propose sampling $m$ labels for each calibration point and using the labels to compute $m$ p-values, then taking the unweighted average of these p-values and applying a transformation to obtain a valid p-variable.
Although Stutz et al. [30] only consider the unweighted average, their transformation is general enough to apply to a weighted average of p-values as well, and can therefore be adapted to our setting.
How the transformations affect the weights
Both our method and the method of Stutz et al. [30] aim to transform a random variable to a p-variable to maintain coverage guarantees. The difference between the two methods lies in the nature of the transformation. Our method applies a linear transformation that preserves the proportions of the weights; this can be important when the weights reflect meaningful quantities, like the weights learned by the routing network of an MoE model. In contrast, Stutz et al. [30] apply a nonlinear, rank-based transformation by computing the empirical CDF of the random variable and returning its value at the observed point. That is, given a random variable $X$, their method estimates its CDF $F$ and uses $F(X)$ as the resulting p-value. While this guarantees validity and preserves ordering, it does not preserve the relative scale between values and therefore discards some of the semantics of the original weights.
Comparing both transformations with our MoE setup
In our method for weighted prediction set aggregation with data-dependent weights, we transform the weighted average of the individual prediction set p-values to a valid p-variable using a linear scaling. This transformation allows us to maintain a coverage guarantee for the prediction set defined by the weighted average of the p-values. In the MoE setting, this prediction set corresponds to aggregating the p-values from each expert according to the weights learned by the routing network.
To adapt the method of Stutz et al. [30] to this context, we substitute their transformation in place of our linear scaling. We now restate their original setting and transformation in more detail to clarify how their method can be extended to our setting.
The transformation proposed by Stutz et al. [30] was originally developed to address the problem of uncertainty in the ground truth labels. In their setting, each calibration point consists only of an input $X_{i}$; they sample $m$ labels to get $m$ p-values for each calibration point, and then take the unweighted average of these p-values. We denote this unweighted average as $P_{\text{avg}}^{i}$, with distribution function $F$. Their method transforms $P_{\text{avg}}^{i}$ into a valid p-value via $F(P_{\text{avg}}^{i})$, and the prediction set is then the set of all labels such that the corrected p-value $F(P_{\text{avg}}^{i})$ corresponding to each sample is greater than some threshold $\alpha$.
Stutz et al. [30] note that the coverage guarantee using the true CDF holds only with an asymptotic number of samples, as the empirical CDF $\widehat{F}$ approaches the true CDF $F$. To establish a finite-sample guarantee, they introduce a DKW-derived correction $\epsilon$ and define their prediction sets based on $\widehat{F}+\epsilon$. We find that, although their proposed prediction set yields an elegant $(1-\alpha)(1-\delta)$ finite-sample guarantee, the $\epsilon$ correction is extremely conservative in practice.
Let us refer to the finite-sample empirical CDF method as ECDF-DKW, and the variant without the DKW correction as simply ECDF. We now present additional experiments where we recreate the main findings of our paper with the ECDF and ECDF-DKW methods.
Figure 9: Regression experiments of our linear transformation method with ECDF and ECDF-DKW. Each row corresponds to a dataset, with plots for marginal coverage, WS coverage, and prediction set size from left to right. ECDF performs similarly (sometimes slightly more conservatively) to WA precise in terms of coverage and efficiency. However, ECDF-DKW is so conservative that it covers the entire label space.
Figure 9 recreates the regression experiments of Figure 2, with the addition of ECDF and ECDF-DKW. We note that ECDF performs very similarly to WA precise in terms of coverage (left), WS coverage (middle), and prediction set size (right), with ECDF being slightly more conservative in most cases. Like WA precise, ECDF is more conservative than split conformal and less conservative than WA targeted, although it still often falls short of the $1-\alpha$ guarantee on the WS slab. On the other hand, ECDF-DKW is so conservative that its prediction sets cover the entire label space. We represent this with 100% coverage on both the coverage plots and WS coverage plots, and we omit ECDF-DKW from the prediction set size plots.
Figure 10: Coverage (top) and prediction set size (bottom) for subgroups with top 50th percentile racial representation in the UCI Communities and Crimes dataset. ECDF performs similarly to WA precise; ECDF-DKW is so conservative that it covers the entire label space. Error bars represent 95% confidence intervals.
Figure 10 recreates the Communities and Crimes experiment of Figure 7 with ECDF and ECDF-DKW. Again, we see that ECDF performs very similarly to WA precise on our demographic-conditioned view of Communities and Crimes, with similar coverage (top) and prediction set size (bottom) across all demographics—with ECDF having slightly lower coverage on most demographics, including demographics where WA precise undercovers. Like before, we also observe that ECDF-DKW is so conservative that it has 100% coverage and unbounded prediction sets.
Why is ECDF-DKW so conservative?
For ECDF-DKW, the finite-sample variant of ECDF, Stutz et al. [30] use DKW to add a finite-sample correction $\epsilon$ to the empirical CDF $\widehat{F}(P_{\text{all}})$, then compare this sum to the significance level $\alpha$. The prediction set with finite-sample guarantees is therefore the set of all labels such that $\widehat{F}(P_{\text{avg}})+\epsilon>\alpha$. However, if $\epsilon$ is already greater than $\alpha$, then this condition is always satisfied and the prediction set includes all labels, becoming unbounded.
The finite-sample correction $\epsilon$ is a function of the number of samples used to compute the empirical CDF. Figure 11 shows that when the merging set size is less than 1000, $\epsilon$ is typically large enough to exceed common values of $\alpha$, making unbounded sets very likely.
Figure 11: Finite-sample correction $\epsilon$ used in ECDF-DKW as a function of the merging set size $|S_{\text{merge}}|$ and user-specified significance level $\delta$. This correction is the offset required to ensure the finite sample guarantee of $(1-\alpha)(1-\delta)$ in Stutz et al. [30]. Dashed lines mark $\alpha$ levels of 0.05 and 0.1. When $\epsilon>\alpha$, the prediction set includes all labels and becomes unbounded, yielding 100% coverage.
Appendix D A note on computational complexity
The computational complexity of our method matches the complexity of existing prediction set aggregation methods for the case of data-independent weights, and includes an additional one-time cost to compute an empirical CDF for the case of data-dependent weights.
The complexity of prediction set aggregation methods was first observed by Cherubin, who compares the computational overhead of their majority vote aggregation method with the overhead of p-value aggregation methods. They note that while majority vote requires simpler operations to determine whether each label is included in the final prediction set, it also requires that predictions be recomputed for each significance level, making it less efficient when sets must be constructed at multiple thresholds. In contrast, p-value methods allow the aggregation to be computed once and then applied to any significance level without additional computation.
These trade-offs in speed and cost may influence which method is better suited to a given application—for example, p-value aggregation may be preferable if prediction sets need to be dynamically thresholded at test time. However, the time complexity of both methods is the same: combining $K$ prediction sets for $N$ test objects with a label space size of $L=|\mathcal{Y}|$ has complexity $\mathcal{O}(KLN)$. This complexity is necessary for all prediction set aggregation methods, as it reflects the cost of evaluating multiple prediction sets across the label space.
With data-independent weights, our method matches this $\mathcal{O}(KLN)$ complexity directly. With data-dependent weights, the only additional computation required is a one-time estimation of a correction factor $\widehat{m}^{*}$ from data split $S_{\text{merge}}$. This step involves computing the empirical CDF of the weighted average p-values on $S_{\text{merge}}$, which has a complexity of $\mathcal{O}(M\log M)$ for a split of size $M$. Importantly, this correction is computed once and does not require retraining, and the rest of the procedure for data-dependent weights has the same $\mathcal{O}(KLN)$ cost as other prediction set aggregation methods.
Appendix E Further implementation details
E.1 Synthetic dataset
For our ablative/expository experiments, we generate a simple homoskedastic dataset to simulate a regression task. Each input is a 16-dimensional vector drawn from a standard normal distribution, and the output label is the sum of the feature values with additive Gaussian noise. Specifically, for each sample $X_{i}$, we have
$y_{i}=\sum_{j=1}^{16}X_{ij}+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ represents additive noise with standard deviation $\sigma=0.1$. We generate both training and test datasets by independently drawing samples and computing the corresponding target labels.
E.2 Mixture-of-Experts model
In all of our experiments, we use an MoE model of $N$ experts, where each expert is a linear model defined as $f_{i}(x;\theta_{i})$. Here, $x$ represents the input features and $\theta_{i}$ the parameters of the $i$th expert. Each expert considers a different subset of features depending on the experiment (§E.3 and §B.1.2). The experts are trained independently using L-BFGS to minimize mean square error (MSE) on the training data.
The routing network is also a linear model, and is responsible for generating a set of weights $\{w_{i}(x)\}_{i=1}^{N}$ that determine the contribution of each expert to the final prediction. To ensure that the weights are positive and sum to one, softmax is applied to the routing network outputs:
$w_{i}(x)=\frac{\exp(g_{i}(x;\phi))}{\sum_{j=1}^{N}\exp(g_{j}(x;\phi))},$
where $g_{i}(x;\phi)$ denotes the output of the routing network for the $i$th expert, and $\phi$ represents the parameters of the routing network.
The routing network is trained to minimize MSE of the aggregated prediction
$\ell_{routing}=\frac{1}{|S_{train}|}\sum_{i\in S_{train}}\left(y_{i}-\sum_{j=1}^{N}w_{j}(x_{i})f_{j}(x_{i};\theta_{j})\right)^{2}.$
In our experiments, the trained MoE model is utilized in two ways. For split conformal, we use the weighted sum of experts as a black-box predictive model. For weighted aggregation, we use the weights from the routing network to scale the p-value functions of each expert.
E.3 Expert feature assignment for UCI experiments
For the UCI experiments, the features for each dataset are partitioned into groups of semantically related features, and each expert in the MoE specializes in a single group of features. We list these groups and their features below, where the feature names are provided with the original dataset [39].
Table 1: Airfoil dataset feature groups
Group Features
Aerodynamics frequency, free-stream-velocity
Geometry attack-angle, chord-length, suction-side-displacement-thickness
Table 2: Wine dataset feature groups
Group Features
Acidity fixed_acidity, volatile_acidity, citric_acid, pH
SugarAndAlcohol residual_sugar, density, alcohol
SulfurAndSalinity chlorides, free_sulfur_dioxide, total_sulfur_dioxide, sulphates
Table 3: Communities dataset feature groups
Group Features
Population population, householdsize, numbUrban, pctUrban, LandArea, PopDens, agePct12t21, agePct12t29, agePct16t24, agePct65up
RaceEthnicity racepctblack, racePctWhite, racePctAsian, racePctHisp, PctForeignBorn, PctSpeakEnglOnly, PctNotSpeakEnglWell, PctBornSameState, PctSameHouse85, PctSameCity85
IncomePoverty medIncome, medFamInc, perCapInc, whitePerCap, blackPerCap, indianPerCap, AsianPerCap, HispPerCap, NumUnderPov, PctPopUnderPov
EmploymentIndustry pctWWage, pctWFarmSelf, pctWInvInc, pctWSocSec, pctWPubAsst, pctWRetire, PctUnemployed, PctEmploy, PctEmplManu, PctEmplProfServ
OccupationEducation PctOccupManu, PctOccupMgmtProf, PctWorkMomYoungKids, PctWorkMom, PctUsePubTrans, PctLess9thGrade, PctNotHSGrad, PctBSorMore, MalePctDivorce, MalePctNevMarr
FamilyStructure FemalePctDiv, TotalPctDiv, PersPerFam, PctFam2Par, PctKids2Par, PctYoungKids2Par, PctTeen2Par, PctLargHouseFam, PctLargHouseOccup, PctSameState85
HousingCharacteristics PersPerOccupHous, PersPerOwnOccHous, PersPerRentOccHous, PctPersOwnOccup, PctHousNoPhone, PctHousLess3BR, MedNumBR, HousVacant, PctHousOccup, PctHousOwnOcc
HousingQualityCosts PctPersDenseHous, PctVacantBoarded, PctVacMore6Mos, MedYrHousBuilt, PctWOFullPlumb, OwnOccLowQuart, OwnOccMedVal, OwnOccHiQuart, RentLowQ, RentMedian
HousingCostsHomelessness RentHighQ, MedRent, MedRentPctHousInc, MedOwnCostPctInc, MedOwnCostPctIncNoMtg, NumInShelters, NumStreet, NumIlleg, PctIlleg, LemasPctOfficDrugUn
Immigration NumImmig, PctImmigRecent, PctImmigRec5, PctImmigRec8, PctImmigRec10, PctRecentImmig, PctRecImmig5, PctRecImmig8, PctRecImmig10
Table 4: Superconductivity dataset feature groups
Feature Group Features
[…]
Table 5: Adult dataset feature groups
Group Features
Demographics age, race, sex
Education education, education-num
Economic Status fnlwgt, capital-gain, capital-loss, hours-per-week
Family Relationship marital-status, relationship
Table 6: Student dataset feature groups
Group Features
Personal Details Marital Status, Gender, Age at enrollment, Nationality, International, Mother’s qualification, Father’s qualification, Mother’s occupation, Father’s occupation
Academic Details Application mode, Application order, Previous qualification, Previous qualification (grade), Admission grade, Daytime/evening attendance, Course, Displaced, Educational special needs
Performance Curricular units 1st sem (credited), Curricular units 1st sem (enrolled), Curricular units 1st sem (evaluations), Curricular units 1st sem (approved), Curricular units 1st sem (grade), Curricular units 1st sem (without evaluations), Curricular units 2nd sem (credited), Curricular units 2nd sem (enrolled), Curricular units 2nd sem (evaluations), Curricular units 2nd sem (approved), Curricular units 2nd sem (grade), Curricular units 2nd sem (without evaluations)
Economic and Financial Scholarship holder, Tuition fees up to date, Debtor, Unemployment rate, Inflation rate, GDP